Calculate the price elasticity of demand when the price of pencils rises from $4.50 to $5.50, when income is $20,000.

1
Expert's answer
2021-10-31T18:29:36-0400

price elasticity of demand

"=\frac{\%\Delta quantity}{\%\Delta price}\\\Delta=\frac{p_2-p_1}{p_1}=\frac{1}{4.5}=0.22\\
\Delta=\frac{Q_2-Q_1}{Q_1}=\frac{-800}{1600}=-0.5\\price \space elasticity\space of \space demand=\frac{-0.5}{0.22}=-2.25"